# A Logarithmic Spiral Formed by a Sequence of Regular Polygons

### v1

@@ 3 Approximating $P_{n}$
$e^{i\frac{\pi}{2}(\gamma+\log 2)}\sum_{k=2}^{n-1}\frac{1}{2}{\left(\frac{k}{\pi}+\frac{k+1}{\pi}-\frac{1}{3}\frac{\pi}{k}-\frac{1}{3}\frac{\pi}{k+1}+\mathcal{O}\left(\frac{1}{k^{3}}\right)\right)}e^{{i\frac{\pi}{2}\left(-\frac{(-1)^{k}}{2k}+\left(\frac{1}{24}+\frac{(-1)^{k}}{4}\right)\frac{1}{k^{2}}+\mathcal{O}\left(\frac{1}{k^{3}}\right)\right)}}$
$e^{i\frac{\pi}{2}(\gamma+\log 2)}\sum_{k=2}^{n-1}{\left(\frac{k+\frac{1}{2}}{\pi}-\frac{\pi}{3}\frac{1}{k+\frac{1}{2}}+\mathcal{O}\left(\frac{1}{k^{2}}\right)\right)}e^{{i\frac{\pi}{2}\left(-\frac{(-1)^{k}}{2}\frac{1}{k+\frac{1}{2}}+\frac{1}{24}\frac{1}{\left(k+\frac{1}{2}\right)^{2}}+\mathcal{O}\left(\frac{1}{k^{3}}\right)\right)}}$
$e^{i\frac{\pi}{2}(\gamma+\log 2)}\sum_{k=2}^{n-1}\left(\frac{k+\frac{1}{2}}{\pi}-\frac{\pi}{3}\frac{1}{k+\frac{1}{2}}+\mathcal{O}\left(\frac{1}{k^{2}}\right)\right)\left(1+i\frac{\pi}{2}\left(-\frac{(-1)^{k}}{2(k+\frac{1}{2})}+\frac{1}{24(k+\frac{1}{2})^{2}}+\mathcal{O}\left(\frac{1}{k^{3}}\right)\right)\right.\\ \left.+\frac{1}{2}\left(i\frac{\pi}{2}\right)^{2}\left(\frac{(-1)^{k}}{2(k+\frac{1}{2})}\right)^{2}+\mathcal{O}\left(\frac{1}{k^{3}}\right)\right)$
@@
$\sim_{\circlearrowleft}\sum_{k=2}^{n-1}{\left(\frac{k+\frac{1}{2}}{\pi}-\frac{\pi}{3}\frac{1}{k+\frac{1}{2}}\right)}\left(1-\frac{i\pi(-1)^{k}}{4}\frac{1}{k+\frac{1}{2}}+\left(\frac{i\pi}{48}-\frac{\pi^{2}}{32}\right)\frac{1}{\left(k+\frac{1}{2}\right)^{2}}\right)$
@@
$\frac{1}{2}\frac{1}{1+i\frac{\pi}{2}}\left(\left(n-\frac{1}{2}\right)^{1+i\frac{\pi}{2}}-\left(n-\frac{3}{2}\right)^{1+i\frac{\pi}{2}}\right) +\frac{1}{2}\left(\left(n-\frac{1}{2}\right)^{i\frac{\pi}{2}}-\left(n-\frac{3}{2}\right)^{i\frac{\pi}{2}}\right).$
Similarly, for $n=2m$,
@@ 4 The Logarithmic Spiral
where $\arg f(t)$ is the continuous branch satisfying $\arg f(t)\sim\frac{\pi}{2}\log t$ as $t\to\infty$.

### v2

@@ 3 Approximating $P_{n}$
$e^{i\frac{\pi}{2}(\gamma+\log 2)}\sum_{k=2}^{n-1}\frac{1}{2}{\left(\frac{k}{\pi}+\frac{k+1}{\pi}-\frac{1}{3}\frac{\pi}{k}-\frac{1}{3}\frac{\pi}{k+1}+\mathcal{O}\left(\frac{1}{k^{3}}\right)\right)}\ e^{{i\frac{\pi}{2}\left(-\frac{(-1)^{k}}{2k}+\left(\frac{1}{24}+\frac{(-1)^{k}}{4}\right)\frac{1}{k^{2}}+\mathcal{O}\left(\frac{1}{k^{3}}\right)\right)}}\left(k+\frac{1}{2}\right)^{i\frac{\pi}{2}}$
$e^{i\frac{\pi}{2}(\gamma+\log 2)}\sum_{k=2}^{n-1}{\left(\frac{k+\frac{1}{2}}{\pi}-\frac{\pi}{3}\frac{1}{k+\frac{1}{2}}+\mathcal{O}\left(\frac{1}{k^{2}}\right)\right)}\ e^{{i\frac{\pi}{2}\left(-\frac{(-1)^{k}}{2}\frac{1}{k+\frac{1}{2}}+\frac{1}{24}\frac{1}{\left(k+\frac{1}{2}\right)^{2}}+\mathcal{O}\left(\frac{1}{k^{3}}\right)\right)}}\left(k+\frac{1}{2}\right)^{i\frac{\pi}{2}}$
$e^{i\frac{\pi}{2}(\gamma+\log 2)}\sum_{k=2}^{n-1}\left(\frac{k+\frac{1}{2}}{\pi}-\frac{\pi}{3}\frac{1}{k+\frac{1}{2}}+\mathcal{O}\left(\frac{1}{k^{2}}\right)\right)\left(1+i\frac{\pi}{2}\left(-\frac{(-1)^{k}}{2(k+\frac{1}{2})}+\frac{1}{24(k+\frac{1}{2})^{2}}+\mathcal{O}\left(\frac{1}{k^{3}}\right)\right)\right.\\ \left.+\frac{1}{2}\left(i\frac{\pi}{2}\right)^{2}\left(\frac{(-1)^{k}}{2(k+\frac{1}{2})}\right)^{2}+\mathcal{O}\left(\frac{1}{k^{3}}\right)\right)\left(k+\frac{1}{2}\right)^{i\frac{\pi}{2}}$
@@
$\sim_{\circlearrowleft}\sum_{k=2}^{n-1}{\left(\frac{k+\frac{1}{2}}{\pi}-\frac{\pi}{3}\frac{1}{k+\frac{1}{2}}\right)}\left(1-\frac{i\pi(-1)^{k}}{4}\frac{1}{k+\frac{1}{2}}+\left(\frac{i\pi}{48}-\frac{\pi^{2}}{32}\right)\frac{1}{\left(k+\frac{1}{2}\right)^{2}}\right)\left(k+\frac{1}{2}\right)^{i\frac{\pi}{2}}$
@@
$\frac{1}{2}\frac{1}{1+i\frac{\pi}{2}}\left(\left(n-\frac{1}{2}\right)^{1+i\frac{\pi}{2}}-\left(n-\frac{3}{2}\right)^{1+i\frac{\pi}{2}}\right) +\frac{1}{2}\left(\left(n-\frac{1}{2}\right)^{i\frac{\pi}{2}}-\left(n-\frac{3}{2}\right)^{i\frac{\pi}{2}}\right)$
(by (3)). Similarly, for $n=2m$,
@@ 4 The Logarithmic Spiral
where $\arg f(t)$ is the continuous branch satisfying $\arg f(t)=\frac{\pi}{2}\log t+o(1)$ as $t\to\infty$.